OpenStudy (elise_a18):

Simlify. square root of 5 over cube root of 5

OpenStudy (january123):

Change these into fractional exponents: the square root of 3 ---------------------------- = the cube root of 2 3^(½) --------- 2^(⅓) Now get a common denominator on both fractional exponents: 3^(3/6) --------- 2^(2/6) To simplify the radical on the bottom we have two 2s so far, but we need 4 more so we can take the 6th root on the bottom. So we will multiply the top and bottom both by 2^(4/6) 3^(3/6)*2^(4/6) ---------------------- = 2^(2/6)*2^(4/6) Exponents on the bottom together: 3^(3/6)*2^(4/6) ---------------------- = 2^(6/6) 3^(3/6)*2^(4/6) ---------------------- = .......2 The top can go back into a radical as the sixth root of these factors: sixth root of [3^(3)*2^(4)] ---------------------------------- = .......2 Since 3^(3)*2^(4) = 27 * 16 = 432, this becomes: sixth root of 432 ----------------------- ............2 That is as simplified as it gets!! I hope that helps!! :-)

OpenStudy (elise_a18):

Here are the options 5 to the power of negative 1 over 6 5 to the power of negative 1 over 6 5 to the power of 5 over 6